What is Ka for the conjugate acid of C5H5N (Kb = 1.7 x 10 9)? 7 3 LO 89 +/- X 100
Answer #1

Ка - Кь - w = (o-19 Ј dissociation constant ka = k, = acid Қаце рах снем к = 1 + xio 9 1. ka = 16 = 169 = 5-88 x166 1. + ҳо